OpenStudy (anonymous):

x*y^2*z^3+x^3*y^2*z+1=x+y+z. Find dz/dx at (1, 1, 1).

OpenStudy (bahrom7893):

what i like doing in these situations is just changing one of the variables to some number, because I know it will be treated as a constant. In this case I would make y some number since we're differentiating z wrt x. So rewrite this as: (by the way any number would do, I just picked 5) x*(5^2)*z^3 + x^3(5^2)z+1 = x + 5 + z It's just something that I prefer doing because it reminds me that I need to treat y as a constant.

OpenStudy (bahrom7893):

so now just use implicit differentiation: (5^2) [x * z^3] + (5^2) [x^3 * z] + 1 = x + 5 + z Whatever's in the brackets is going to use the product rule.

OpenStudy (bahrom7893):

I just decided to change the 5s back to ys because I somewhat confused myself. (y^2) [ x * 3z^2 (dz/dx) + z^3 ] + (y^2) [x^3 * (dz/dx) + z * 3x^2] = 1 + 0 + dz/dx

OpenStudy (bahrom7893):

The 0 on the right hand side is coming from differentiating y - as far as dz/dx is concerned, y is a constant, it does not involve any z or x terms

OpenStudy (bahrom7893):

So now just do the algebra and get dz/dx on its own.

OpenStudy (anonymous):

Wait a minute, let me do the work.

OpenStudy (anonymous):

Thank you so much, I've got the correct answer.

OpenStudy (bahrom7893):

Good job